OpenStudy (anonymous):

limit as x approaches quantity pi divided by two of quantity three times e to the x times cosine of x. WILL MEDAL!!

OpenStudy (anonymous):

@IrishBoy123 Can you help?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}(3e^xcosx)\]

OpenStudy (irishboy123):

quicky, have you plugged in \(x = \pi/2\)? i will latex what i think you are looking at

OpenStudy (anonymous):

Yes when plugged in the cos(pi/2) = 0 and e^0 = 1, so would the limit then only be 3? or would I multiply 3 x 0 making the limit 0?

OpenStudy (irishboy123):

\(\large lim_{x \rightarrow \ \frac{pi}{2}} \ \ 3e^{x \ cosx}\) ???

OpenStudy (irishboy123):

if so the exponent is \(\pi/2 * 0 = 0\)

OpenStudy (anonymous):

I have the equation written out above, the cos x isn't part of the exponent.

OpenStudy (irishboy123):

oh sorry, will look at yours

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}(3e^2cosx)\] @IrishBoy123

OpenStudy (irishboy123):

\(3 e^{\pi/2} \times 0 = 0\)

OpenStudy (anonymous):

Oh ok so don't worry about e^pi/2 because the x0 from the cos(pi/2) already makes the entire limit 0?

OpenStudy (irishboy123):

yes

OpenStudy (anonymous):

Ok thank you so much! You've been a huge help. (:

OpenStudy (irishboy123):

finally, is this it?!?! \[\lim_{x \rightarrow \frac{\pi}{2} } \ 3e^x \ \cos x\]

OpenStudy (irishboy123):

just to be clear...because that will be zero